### v2

# Floorplan-SLAM: A Real-Time, High-Accuracy, and Long-Term Multi-Session Point-Plane SLAM for Efficient Floorplan Reconstruction

Abstract
[…]
I INTRODUCTION
Floorplans are crucial for indoor robot navigation and high-level scene understanding, and are widely applied in scenarios ranging from service robots in assisted living environments to autonomous drones operating in complex indoor facilities. In these scenarios, generating a floorplan in real-time is particularly advantageous, as it not only provides geometric structure for accurate localization but also furnishes semantic cues for higher-level decision-making. Despite the growing need for real-time floorplan reconstruction, existing approaches often present significant limitations. Some methods operate in an offline manner, requiring a complete map of the environment prior to reconstruction [1, 2, 3]. Others rely on expensive sensors such as LiDAR or RGB-D cameras [4, 5], or computationally intensive neural networks [6, 7, 8], which increase both deployment costs and system complexity. While monocular or stereo camera-based SLAM has attempted to use high-level features such as planes for real-time vectorized map construction [9, 10, 11], these methods lack high-level scene layout understanding, thereby failing to achieve complete and compact indoor vectorized reconstruction. Additionally, they cannot perform robust plane extraction in weakly textured scenes.
Figure 1: Floorplan reconstruction results on the meeting rooms sequence from the self-collected dataset by Floorplan-SLAM. Floorplan-SLAM can robustly process stereo images containing low-texture and cluttered objects at 45 FPS on a CPU, and it merges multiple sub-sessions recorded at different times, ultimately achieving real-time, high-accuracy, and long-term floorplan reconstruction.
To address the above problems, an effective real-time solution that works with more accessible and cost-effective sensors, such as stereo cameras, remains uncertain. In this work, we present Floorplan-SLAM, a novel framework that tightly integrates floorplan reconstruction into a multi-session SLAM system by seamlessly interacting with plane extraction, pose estimation, back-end optimization, and loop & map merging. Unlike prior methods, our approach operates exclusively with stereo cameras under standard computing resources, allowing real-time performance in large-scale indoor settings. A robust plane extraction algorithm is first developed to handle weakly textured surfaces by leveraging two spatially complementary features in a compact plane parameter space. We further introduce a floorplan reconstruction module tightly integrated with the SLAM system that formulates and solves an innovative optimization problem using the continuously optimized plane landmarks and poses from the SLAM system, enabling real-time incremental floorplan reconstruction. Furthermore, by leveraging the map merging capability of multi-session SLAM, our method supports long-term floorplan reconstruction across multiple sessions without redundant data collection, making it more efficient for large-scale environments. To the best of the authors’ knowledge, this is the first work that relies solely on a stereo camera and standard CPUs to achieve real-time floorplan reconstruction in large-scale indoor environments, making it well-suited for various robotics applications with strict cost and real-time requirements.
The contributions of our work can be summarized as follows:
  1. 1.
    We propose a robust plane extraction algorithm that operates in a compact plane parameter space and leverages two spatially complementary features to accurately and completely detect planar structures, even in weakly textured scenes.
  2. 2.
    We introduce a floorplan reconstruction module that is tightly coupled with the multi-session SLAM system, leveraging continuously optimized plane landmarks and poses from the SLAM system to formulate and solve a novel binary linear programming problem with trajectory constraint thereby enabling real-time, high-accuracy, and long-term floorplan reconstruction.
  3. 3.
    Experiments on the VECtor and our self-collected datasets demonstrate that Floorplan-SLAM significantly outperforms state-of-the-art methods in terms of plane extraction robustness, pose estimation accuracy, and floorplan reconstruction fidelity and speed, achieving real-time performance at 25–45 FPS without GPU acceleration.
II Related Work
Floorplan reconstruction aims to convert raw sensor data into vectorized geometric models. Some primitives-based vectorized reconstruction methods [1, 2, 3] detect planes from the input point cloud and then perform a global optimization to select the optimal subset of candidate planes for vectorized reconstruction. However, these methods usually require a prior global map, such as a dense point cloud or a mesh map of the entire scene, which is difficult to obtain from actual point clouds or reconstructed dense meshes due to noise, outliers, and missing data.
With the development of neural networks, deep learning-based floorplan reconstruction methods have become of interest. These methods [6, 7, 8] typically first project the input dense point cloud onto a top view to create a 2D density map, and then use neural networks to infer the floorplan of the scene from the projected density map. However, these methods are only applicable to specific scenarios, exhibit limited generalizability, and incur high time cost and computational resource consumption.
To achieve real-time floorplan reconstruction, some works [9, 4, 5] utilized visual SLAM to extract plane features from image sequences, which are then used to create landmarks for map construction. However, these methods either lack high-level scene layout understanding [9], thereby failing to achieve complete and compact indoor vectorized reconstruction, or integrate the acquired poses and scene structures in a loosely coupled manner [4, 5], leading to poor system robustness due to the inability to perform mutual optimization between pose and scene structure. In addition, these methods can only process one set of consecutive images at a time, making it impossible to merge multiple sequences and achieve long-term mapping.
Furthermore, these methods typically utilize RGB-D [5] or neural networks [9, 4] for plane extraction, which is challenging for stereo cameras. In [11], a plane extraction method based on intersecting lines that satisfy specific geometric constraints was proposed. However, this method cannot accurately distinguish real planes from pseudo-planes and is highly susceptible to noise. To enhance the robustness of plane extraction, a region-growing-based plane extraction method was introduced in [10]. Nevertheless, it struggles to extract complete planes in weakly textured scenes and does not fully leverage visual information.
Figure 2: System Overview. Floorplan-SLAM tightly integrates a stereo point-plane SLAM system with a floorplan reconstruction module. The former provides accurate plane landmarks and trajectory information, while the latter formulates and solves a novel optimization problem, ultimately achieving real-time, high-accuracy, and long-term floorplan reconstruction.
III Point-Plane-Based Stereo SLAM System
In this section, we introduce the proposed point-plane-based stereo SLAM system. The system first uses two sets of spatially complementary point features to extract planes, and then employs both point and plane features to estimate poses. Plane landmarks are subsequently created and managed for floorplan reconstruction. In addition, a multi-session map merging mechanism is introduced to support long-term mapping. The overview of the system is illustrated on the left side of Fig. 2. We refer the reader to [10] for more details on the pose estimation. Next, we focus on describing the operations related to plane extraction, plane landmark management, and multi-session map merging.
III-A Plane Extraction
III-A1 Support Point Extraction
We employ both the Sobel and ORB features for plane extraction, based on the insight that these two types of features focus on distinct yet complementary regions of a plane: Sobel predominantly captures boundary edges, while ORB detects interior corner points. By harnessing their complementary strengths, we achieve more robust and comprehensive plane extraction. Specifically, we first rectify the input stereo pair, ensuring that stereo correspondences are restricted to the same row in both images. Next, we perform ORB and Sobel feature matching across the stereo images to establish a set of correspondences, followed by consistency checks and outlier removal, thus obtaining a collection of 2D support points. For a 2D support point with pixel coordinates $\left(u,v\right)$ and disparity value $d$, its 3D coordinates $\mathbf{p}_{s}$ are computed through triangulation as follows:
$\mathbf{p}_{s}=\left[\begin{matrix}x&y&z\\ \end{matrix}\right]^{\top}=\left[\begin{matrix}\frac{\left(u-c_{x}\right)z}{f_{x}}&\frac{\left(v-c_{y}\right)z}{f_{y}}&\frac{f_{x}b}{d}\\ \end{matrix}\right]^{\top},$ (1)
where $\left(f_{x},f_{y}\right)$ is the focal length, $\left(c_{x},c_{y}\right)$ is the principal point, and $b$ is the baseline, all known from calibration.
III-A2 Mesh Generation
Because the support points acquired through stereo matching are relatively sparse, the estimated normal vectors of these points exhibit significant errors, rendering them unsuitable as reliable information for plane extraction. Therefore, we employ a set of triangles obtained via Delaunay triangulation [12] on the 2D support points, with these support points serving as vertices, as the fundamental units for plane extraction. Then, we project these triangles into 3D space using the 3D coordinates of the support points to obtain a 3D mesh. We further prune the mesh by discarding triangles in the 3D mesh with long edges, high aspect ratios, and small acute angles [13].
III-A3 Plane Parameter Space Construction
Each triangle in the 3D mesh is considered a planar patch, represented by $\pi=\left[\mathbf{n}^{\top},d\right]^{\top}$ in Cartesian space, where $\mathbf{n}=\left[n_{x},n_{y},n_{z}\right]^{\top}$ is the unit normal vector calculated by the cross product of two triangle edges, and $d$ is the distance from the origin to this triangle planar patch. We project these triangle planar patches from Cartesian space into Plane Parameter Space (PPS) [14], which compactly represents planes in Cartesian space. Specifically, given a triangle planar patch in Cartesian space, its representation in the PPS is a point $\mathbf{p}_{\pi}$:
$\mathbf{p}_{\boldsymbol{\pi}}=\left[\begin{array}[]{l}\phi\\ \psi\\ d\\ \end{array}\right]=\left[\begin{array}[]{c}\mathrm{atan2}\left(n_{y},n_{x}\right)\\ \mathrm{arccos}\left(n_{z}\right)\\ d\\ \end{array}\right],$ (2)
where $\phi$ and $\psi$ are the azimuth and elevation angles of the normal vector respectively.
III-A4 Plane Parameters Estimation
Ideally, the coplanar triangle planar patches in Cartesian space share the same coordinates in the PPS. However, due to noise, the coordinates of these coplanar triangle planar patches in the PPS are not exactly the same but rather close. Therefore, the problem of detecting coplanar triangle planar patches in Cartesian space is transformed into a point clustering problem in the PPS. We employ DBSCAN [15] to cluster the points in the PPS, which is robust to noise.
After obtaining the coplanar triangle planar patches, the vertices of these triangles are coplanar points. We apply RANSAC [16] to these coplanar points to obtain accurate plane parameters, and only planes with an inlier ratio exceeding $\theta_{i}$ are added to the plane feature set. In our experiments, $\theta_{i}$ is set to 0.75.
III-B Plane Landmark Management
For any two plane landmarks $\mathbf{\Pi}_{i}$ and $\mathbf{\Pi}_{j}$, if the angle between their normal vectors is less than $\theta_{c}$ and the minimum point-to-plane distances from the support points of $\mathbf{\Pi}_{i}$ to $\mathbf{\Pi}_{j}$ and from $\mathbf{\Pi}_{j}$ to $\mathbf{\Pi}_{i}$ are both less than $d_{c}$, we remove the plane landmark with fewer keyframe observations and transfer its observation information and support points to another plane landmark, followed by RANSAC to update the plane points. The above operations continue throughout the entire process of the system. In our experiments, $d_{a}$ is set to 1 cm, $\theta_{c}$ is set to 5, and $d_{c}$ is set to 2 cm.
III-C Multi-Session Map Merging
We use DBoW2 [17] for place recognition and compute the aligning transformation $\mathbf{T}_{cm}$ between the two maps. Then, a local window is defined, comprising matching keyframes from both maps, their respective neighbors in the covisibility graph, and the observed point and plane landmarks. Duplicated landmarks from the current map are removed within this local window, and their observations are transferred to the corresponding landmarks in the matching map. A local bundle adjustment (BA) is performed to optimize all keyframes and the point and plane landmarks they observe within the local window. Pose graph optimization is then conducted on the essential graph of the merged map, with the keyframes in the local window fixed. Finally, a global BA is executed in a separate thread to further optimize the merged map and enhance overall consistency. By merging sequences captured at different times, we do not need to capture all map areas in a single pass, which is crucial for large-scale scene long-term mapping, as it significantly reduces the computational and storage burden on the acquisition devices. Notably, during each BA process, the parameters of plane landmark $\mathbf{\Pi}_{k}$ are optimized, and its support points are updated by aggregating the support points of its observations $\boldsymbol{\pi}_{i,k}$ in the keyframes into $\mathbf{\Pi}_{k}$ followed by RANSAC, thereby providing a more accurate scene layout for the subsequent floorplan reconstruction.
IV Floorplan Reconstruction
The floorplan reconstruction module runs in parallel with the tracking, local mapping, and loop & map merging threads. It processes the plane landmarks and trajectory information generated by the SLAM system and reconstructs the floorplan by formulating and solving a novel optimization problem. An overview of this workflow is shown on the right side of Fig. 2.
IV-A Candidate Wall Segment Generation
We first select valid plane landmarks that are approximately perpendicular to the ground. We then project their plane points onto the ground to obtain projection lines and support points, thereby generating 2D wall segments. We further merge segments whose angle is less than $\theta_{r}$ and share more than $n_{r}$ support points sufficiently close to both segments, in order to improve the regularity of the wall segments. Next, we extend and pairwise intersect these segments to generate candidate wall segments. In our experiments, $\theta_{r}$ is set to 10, and $n_{r}=\min\left(\left|\mathrm{SP}\left(s_{i}\right)\right|,\left|\mathrm{SP}\left(s_{j}\right)\right|\right)/10$, where $\left|\mathrm{SP}\left(s_{i}\right)\right|$ represents the number of support points of $s_{i}$.
IV-B Wall Segment Selection
Given a set of candidate wall segments $C=\left\{c_{1},c_{2},\dots,c_{N}\right\}$ introduced by the pairwise intersection process, we aim to select an optimal subset of this set through global optimization, thereby achieving accurate floorplan reconstruction. We formulate the optimization problem as a binary linear programming problem under hard constraints and solve it via an energy minimization approach. The objective function of this optimization problem consists of three energy terms, the point fitting term $E_{f}$, the point coverage term $E_{c}$, and the model complexity term $E_{m}$, subject to multiple hard constraints.
IV-B1 Point Fitting Term
This term aims to evaluate the fitting degree between candidate wall segments and their support point sets:
$E_{f}=1-\frac{1}{\left|P\right|}\sum_{i=1}^{N}{f\left(c_{i}\right)x_{i}},$ (3)
$f\left(c\right)=\sum_{p\in\mathrm{SP}\left(c\right)\left|\mathrm{dist}\left(p,c\right)<\varepsilon_{f}\right.}{\left(1-\frac{\mathrm{dist}\left(p,c\right)}{\varepsilon_{f}}\right)},$ (4)
where $\left|P\right|$ is the total number of support points of candidate wall segments, $\mathrm{SP}\left(c\right)$ is the support point set of segment $c$, and $\mathrm{dist}\left(p,c\right)$ is the distance from point $p$ to segment $c$. Only points with a distance less than $\varepsilon_{f}$ to the corresponding segment are used in the calculation of $f\left(c\right)$, and $\varepsilon_{f}$ is set to the average distance between candidate wall segments and their support point sets in our experiment.
IV-B2 Point Coverage Term
This term aims to appropriately evaluate the coverage degree between candidate wall segments and their support point sets in cases where the point cloud is incomplete due to missing data. We consider a segment to be covered by two adjacent support points if the distance between these points is less than $\varepsilon_{c}$:
$E_{c}=\frac{1}{N}\sum_{i=1}^{N}{\left(1-\frac{\mathrm{len}_{\mathrm{cov}}\left(c_{i}\right)}{\mathrm{len}\left(c_{i}\right)}\right)x_{i}},$ (5)
where $\mathrm{len}_{\mathrm{cov}}\left(c_{i}\right)$ is the total length of the portion of $c_{i}$ covered by its support points, and $\mathrm{len}\left(c_{i}\right)$ is the length of $c_{i}$. In our experiments, $\varepsilon_{c}=10\cdot\mathrm{density}\left(P\right)$, where $\mathrm{density}\left(P\right)$ represents the average of the mean distances between all support points and their respective 10-nearest neighbors.
IV-B3 Model Complexity Term
This term aims to evaluate the complexity of the reconstructed floorplan and maintain it at an appropriate level of complexity. Specifically, if an intersection point $v_{i}$ generated by pairwise intersection of the wall segments is connected by non-collinear candidate wall segments $c_{i}$ and $c_{j}$, and both $c_{i}$ and $c_{j}$ are selected, we consider $v_{i}$ to introduce a sharp structure:
$E_{m}=\frac{1}{M}\sum_{i=1}^{M}{\mathbb{I}_{\mathrm{sharp}}\left(v_{i}\right)},$ (6)
where $\mathbb{I}_{\mathrm{sharp}}\left(v_{i}\right)$ is an indicator function that equals 1 if $v_{i}$ introduces a sharp structure, and 0 otherwise.
Figure 3: Comparison of floorplan reconstruction results with and without trajectory constraints. The top-left shows the CAD floor plans provided by the meeting room scenario in the self-collected dataset. The red box indicates the reconstruction results without trajectory constraints, while the green box represents the reconstruction results with trajectory constraints. The trajectory constraints ensure that the reconstructed floorplan contains navigable openings, thereby significantly improving reconstruction accuracy.
IV-B4 Constraints
Trajectory information is a unique component of the SLAM system, as it inherently contains prior knowledge about navigable areas, as shown in Fig. 3. We exclude candidate wall segments that are crossed by the trajectory from the selection:
$\sum_{i=1}^{N}\mathbb{I}_{\mathrm{cross}}\left(c_{i}\right)\cdot x_{i}=0,$ (7)
where $\mathbb{I}_{\mathrm{cross}}\left(c_{i}\right)$ is an indicator function that equals 1 if $c_{i}$ is crossed by the trajectory, and 0 otherwise. To ensure that the floorplan is closed except at the passable areas, the intersection point $v_{i}$, which is not connected by segments crossed by the trajectory, must be connected by at least two and at most four candidate wall segments. Otherwise, it must be connected by at least one and at most $4-n_{c}$ candidate wall segments:
$\sum_{i=1}^{N}{x_{i}}&\in\left\{0,2,3,4\right\},&\text{if}\,\,\forall c_{i}\in\mathcal{C}\left(v_{j}\right),\,\mathbb{I}\left(c_{i}\right)=0,\\ 0\leq\sum_{i=1}^{N}{x_{i}}&\leq 4-n_{c},&\text{if}\,\,\exists c_{i}\in\mathcal{C}\left(v_{j}\right),\,\mathbb{I}\left(c_{i}\right)=1,\end{aligned}\right.$ (8)
where $\mathcal{C}\left(v_{j}\right)$ and $n_{c}$ represent the set of candidate wall segments connected to $v_{j}$ and the number of segments in the set that are traversed by the trajectory, respectively.
IV-B5 Optimization
The objective function $E$ of this optimization problem is the weighted sum of the above energy terms:
$E=\lambda_{f}E_{f}+\lambda_{c}E_{c}+\lambda_{m}E_{m}.$ (9)
We utilize the SCIP solver [18] to minimize (9) subject to the hard constraints (7) and (8) to obtain the optimal subset of candidate wall segments. Then, the floorplan is reconstructed by assembling the segments in the optimal subset.
V Experiment
V-A Plane Extraction Performance Evaluation
Since we utilize the extracted planes for floorplan reconstruction, the accuracy and stability of these planes are crucial to the overall quality of the reconstruction outcomes. We evaluate our plane extraction algorithm and compare it with the plane extraction modules of two point-plane-based stereo SLAM systems, Stereo-Plane-SLAM [11] and RSS [10]. We use plane reprojection error and plane observation count as evaluation metrics to measure the accuracy and stability of the extracted planes, respectively. Specifically, the plane reprojection error represents the matching error between plane observations and landmarks, while the plane observation count indicates the average number of times each valid plane landmark is observed by keyframes. To eliminate the impact of trajectory errors on plane extraction performance, we utilize ground-truth trajectories for evaluation.
Table I compares the plane extraction performance of different methods on the VECtor and self-collected datasets. Our method significantly outperforms SP-SLAM and RSS in both accuracy and stability. Because RSS uses a region-growing strategy for plane extraction, insufficient support points in weakly textured regions can lead to multiple planes being extracted from a single surface, making them more prone to noise and thus less stable. In contrast, our method employs DBSCAN clustering in the PPS space to robustly detect complete and accurate discontinuous planes even in weakly textured scenes lacking support points. Additionally, by using both ORB and Sobel features, which focus on corner and edge regions respectively, our method can extract more planes than RSS, which relies solely on Sobel features, particularly in VECtor scenarios with numerous patches on ceilings and floors, as illustrated in Fig. 4. SP-SLAM relies on intersecting lines for plane extraction. Although many line segments can be identified, strict filtering to avoid spurious planes leaves only a few that contribute to plane extraction. Moreover, these segments are highly sensitive to noise, making the extracted planes unstable across multiple frames and reducing their overall accuracy and stability.
TABLE I: Plane Extraction Performance Comparison on the VECtor and Self-Collected Datasets
Sequence Reprojection Error $\downarrow$ Observation Count $\uparrow$
SP1 RSS Ours SP1 RSS Ours
corridors 0.033 0.028 0.009 11.909 32.059 103.727
school 0.039 0.029 0.011 6.250 28.265 79.375
meeting rooms 0.048 0.017 0.007 15.333 34.867 226.286
offices 0.062 0.022 0.008 19.750 22.528 135.167
café 0.057 0.024 0.008 14.333 20.286 147.250
  • 1
    SP is the abbreviation for Stereo-Plane-SLAM [11].
Figure 4: Comparison of plane extraction results on the VECtor (upper) and self-collected (lower) datasets. ORB features are represented by crosses, and Sobel features are represented by dots in our method. The plane boundaries in both RSS and our approach are determined by the convex hull of the corresponding plane points.
V-B Localization Accuracy Evaluation
Figure 5: Our self-collected dataset setup.
Accurate poses allow planes extracted from different frames to be seamlessly integrated into the global map, maintaining the global consistency of the incrementally reconstructed floorplan. We evaluate the localization performance of our method in large-scale indoor scenes and compare it with the most relevant state-of-the-art visual SLAM systems, including ORB-SLAM3 (stereo mode) [19], Stereo-Plane-SLAM [11], and RSS [10]. To assess the impact of plane features, we also conduct an ablation study using a point-only variant of our system. Experiments are performed on the public VECtor dataset [20] and a self-collected dataset, both containing sequences exceeding 100 meters. The self-collected dataset was captured using a ZED2 camera with GeoSLAM ZEB as ground truth (Fig. 5). We align the estimated trajectory with ground truth using the Umeyama algorithm without scaling and use the RMSE of the absolute trajectory error (ATE) to measure global drift. Each sequence is run ten times to mitigate multi-thread randomness, and the median results are reported.
TABLE II: Localization Performance Comparison on the VECtor and Self-Collected Dataset (ATE RMSE [m])
Sequence ORB1 SP1 RSS PO1 Ours
corridors_dolly 0.96 0.95 0.94 0.96 0.93
units_dolly 2.48 2.49 1.81 2.42 1.64
units_scooter 1.89 1.82 1.74 1.84 1.41
school_dolly 1.40 1.39 1.40 1.39 1.37
school_scooter 1.39 1.30 1.33 1.38 1.31
Corridor 1.53 1.57 1.38 1.54 1.29
Room 1.40 1.25 1.18 1.38 1.07
Stair 2.35 1.79 1.66 2.40 1.60
Basement 2.49 2.41 1.61 2.52 1.56
  • 1
    ORB, SP, and PO are abbreviations for ORB-SLAM3, Stereo-Plane-SLAM, and the point-only variant of our system, respectively.
As shown in Table II, our method outperforms ORB-SLAM3 and the Point-Only variant across all sequences, especially on the units sequences of the VECtor dataset and the Basement sequence of the self-collected dataset, which feature challenging scenes characterized by low texture, dynamically changing illumination, and motion blur. In these scenarios, our approach achieves a notable accuracy improvement of 0.5 – 0.9 m. This can be attributed to two primary reasons: (1) Compared with point correspondences, plane correspondences are more accurate and stable, particularly in the aforementioned challenging scenarios; (2) Planar structures, especially life-long structures such as walls with extensive spatial coverage, can be tracked across considerable distances, thereby providing comprehensive and reliable constraints for pose estimation. Furthermore, in comparison with the two point-plane-based stereo SLAM systems, our method yields the smallest average RMSE and still demonstrates superior performance on these challenging sequences. This can be attributed to the more comprehensive and accurate plane constraints introduced by our advanced plane extraction method, as demonstrated in Section V-A.
Figure 6: Floorplan reconstruction results on the VECtor dataset of RoomFormer, VecIM, and Ours. We compute and visualize the mean and RMS Hausdorff error (m) between the reconstructed models and the LiDAR point cloud ground truth. The error distribution is color-coded, where blue indicates smaller errors and red represents larger errors.
V-C Floorplan Reconstruction Performance Evaluation
We evaluate the floorplan reconstruction performance of our method on the VECtor [1] and a self-collected dataset, comparing it with two state-of-the-art offline floorplan reconstruction methods: VecIM [1] and RoomFormer [6]. The former employs a gravity-aligned point cloud as input, while the latter uses a 2D density map obtained by projecting the point cloud along the gravity axis. In contrast, our method uses only stereo images as input. VecIM and RoomFormer both utilize a dense multi-view stereo (MVS) point cloud reconstructed by COLMAP [21], which takes advantage of more comprehensive global optimization and additional multi-view information. This results in a more accurate point cloud than that generated by our SLAM system, ensuring fairness in the comparison. We uniformly sample 10K points from the reconstructed floorplan and compute the Hausdorff distance between these sampled points and the ground-truth point cloud to evaluate the reconstruction accuracy, using a dense, high-precision LiDAR point cloud as the ground truth.
V-C1 Evaluation on the VECtor Dataset
Figure 6 shows the qualitative and quantitative results on the VECtor dataset, where our method achieves the lowest mean Hausdorff error and reconstructs a more regular floorplan closer to the real scene. Since VecIM directly applies efficient RANSAC to detect planes from the entire point cloud, it is highly susceptible to outliers and noise, which is an inevitable issue for MVS point clouds in weakly textured scenes. As a result, it extracts many incorrect planes, leading to numerous redundant layered structures. Moreover, VecIM enforces the floorplan to be closed, resulting in many erroneous closed regions. In contrast, although our method also faces challenges in weakly textured regions, it leverages plane points from reliable plane landmarks that persist across multiple frames and undergo continuous optimization. While these points are sparser, they are robust matches carefully selected through strict consistency checks and outlier removal, resulting in significantly higher accuracy and structural regularity. Moreover, by incorporating trajectory information, our method correctly handles navigable areas and openings, avoiding erroneous closed regions.
Figure 7: Floorplan reconstruction results on the self-collected dataset of VecIM, RoomFormer and Ours. For each reconstruction result, the mean Hausdorff error (m) and the error model are computed from the output to the ground truth. From top to bottom, we display the office, café, and meeting rooms, respectively. Our method achieves the best performance in both reconstruction accuracy and regularity.
V-C2 Evaluation on the Self-collected Dataset
To evaluate the floorplan reconstruction performance of our method in complex indoor environments containing numerous non-lifelong structures (e.g., tables, chairs), and to assess its long-term floorplan reconstruction capability, we used a ZED2 camera to collect data in three large-scale scenes, covering meeting rooms, offices, and a café, among other places, each covering an area of 1000 square meters. In the meeting rooms scenario, due to the inconsistent accessibility of different conference rooms, we conducted five separate data collections over two days, with each sub-session covering a different region of the scene, as shown in Fig. 1. The qualitative and quantitative results of floorplan reconstruction on the self-collected dataset are presented in Fig. 7. Our method still reconstructs the most regular and faithful floorplan. Benefiting from rich textures, the MVS point cloud exhibits relatively low error, allowing VecIM to accurately reconstruct main structures. However, VecIM is unable to focus solely on extracting lifelong structures and detects numerous temporary structures such as table sides, adding redundant elements and failing to handle navigable areas effectively. In contrast, although our method may initially detect planes from non-lifelong objects, these planes typically do not persist across multiple frames and are not considered valid plane landmarks for the floorplan reconstruction module, enabling a focus on lifelong structures. Moreover, by fully leveraging trajectory information, we effectively handle openings in navigable areas, such as corridors and room entrances, ensuring that the reconstructed floorplan accurately represents passable regions. In addition, although our self-collected dataset contains room scenes, RoomFormer lacks generalizability for large-scale environments and thus can only generate a limited number of correctly identified rooms. Notably, by sequentially merging multiple subsequences, our method yields a complete large-scale floorplan that encompasses multiple meeting rooms, exhibiting remarkable accuracy and regularity, which strongly demonstrates the long-term globally consistent mapping capability of our approach. As a result, there is no need to acquire the complete scene data in a single pass, greatly reducing the complexity of data collection and reconstruction.
V-D Runtime Analysis
TABLE III: Execution time comparison (in milliseconds).
Set. Sys. ORB-SLAM3 Ours
Seq.
school
1224$\times$1024
Meeting
640$\times$360
school
1224$\times$1024
Meeting
640$\times$360
Tra. PE1 28.77 13.58 28.96 13.51
PlE1 - - 29.92 14.01
PP1 8.44 5.54 11.2 8.32
Total 37.36 19.28 41.38 22.61
Rec. SG1 - - 42.30 37.37
SS1 - - 39.32 55.50
Total - - 81.62 92.87
  • 1
    PE, PlE, PP, SG, and SS are the abbreviations for ORB Point Feature Extraction, Plane Feature Extraction, Pose Prediction, Segment Generation, and Segment Selection, respectively.
The runtime analysis was conducted on a desktop with an AMD Ryzen 7 4800H CPU (running at 2.9 GHz) and 32 GB of RAM. We evaluated our method on the school sequence from the VECtor dataset and on the meeting rooms sequence from our self-collected dataset, comparing the results with ORB-SLAM3. Table III shows the runtimes of the two threads where our method and ORB-SLAM3 differ most; the other threads show comparable execution times. We extract Sobel and ORB features in parallel CPU threads, with the former being faster. Moreover, ORB extraction is the main time-consuming step compared to other plane extraction operations, such as mesh generation, clustering, and plane parameter estimation. As a result, our overall plane extraction time is close to that of ORB feature extraction. The floorplan reconstruction module comprises two primary components: segment generation and selection. We evaluate the runtime by measuring the duration of the final reconstruction in each sequence, which is the most time-consuming due to the largest number of segments. Notably, considering that valid plane landmarks are typically updated every several frames, we perform floorplan reconstruction every five frames. This frequency is sufficient to provide timely updates, thereby enabling real-time floorplan reconstruction. We also recorded the total reconstruction time consumed by VecIM, including the MVS dense point cloud reconstruction and the floorplan reconstruction stages. The first stage took 16 hours and 38.9 minutes, and the second stage took 5.1 minutes, resulting in a total of 16 hours and 44 minutes. In contrast, we processed only stereo images at 45 FPS, and completed the entire floorplan reconstruction in 9.4 minutes, which is two orders of magnitude faster than VecIM. This result strongly demonstrates the efficiency and practicality of our method.
VI Conclusion
[…]